Show that if n|m, where n and m are integers greater than 1, and if a≡b(mod m), where a and b are integers, then a≡b(mod n).
1
Expert's answer
2020-10-20T18:25:45-0400

"\mathbf{Given:\;n|m\;where\;n\;and \;m\;are\;integers\;greater\;than\;1,and}"

"\mathbf{a\equiv b(mod\;m),\;where\;a\;and\;b\;are\;integers.}"


"\mathbf{To\;show: a\equiv b(mod\;n)}"


"\mathbf{Proof:\;Since\;n|m,i.e,\;n\;divides\;m,we\;have-}"

"\mathbf{\;\;\;\;\;\;\;\;\;\;\;\;\;\;m=n\alpha,where\;\alpha\;is\;an\;integer.\;\;\;\;................(A)}"


"\mathbf{Also\;a\equiv b(mod\;m) \implies m\;divides\;(a-b)\implies m|(a-b)}"

"\mathbf{\;\;\;\implies a-b=m\lambda,where\;\lambda\;is\;an\;integer.\;\;\;................(B)}"


"\mathbf{Substitute\;m=n\alpha\;in\;equation\;(B),we\;get-}"

"\mathbf{a-b=(n\alpha)\lambda}"

"\implies\mathbf{a-b=n(\alpha\lambda)\;\;\;\;\;(\because Multiplication\;is\;associative.)}"

"\implies \mathbf{n\;divides\;a-b}"

"\implies \mathbf{n|(a-b)}"

"\implies \mathbf{a-b\equiv0(mod\;n)}"

"\implies \mathbf{a\equiv b(mod\;n)}"

"\mathbf{Hence\;the\;result.}"